# Emerging Technologies for 6G Non-Terrestrial-Networks: From Academia to Industrial Applications

### v1

@@ III-B3 Routing
III-C Space-Air-Ground Integrated Networks
Space-Air-Ground Integrated Networks (SAGINs) aim to achieve full network coverage and ubiquitous services by integrating terrestrial networks, satellite networks and aerial networks[70]. While applying network slicing to SAGINs is an effective solution for efficient usage of network resources, this technique faces new challenges such as complex slice orchestration in multi-domain networks, resource optimization considering UAVs position, and dispatching cost. To address those challenges, multiple innovative schemes have been proposed.
III-C1 Resource Allocation
For instance, as the main concern of network slicing application, resource allocation is discussed in [71, 72, 73, 74]. Particularly, in [71], the authors propose a framework for integrating network slicing to a SAGIN, which establishes three types of RAN slices, i.e., high-throughput, low-delay, and wide-coverage. In particular, a non-scalar multi-objective optimization problem (MOOP) is formulated to jointly optimize throughput, service delay and coverage area. Moreover, a Central and Distributed Multi-agent Deep Deterministic Policy Gradient (CDMADDPG) algorithm is developed to solve the problem. This CDMADDPG algorithm uses a centralized unit to determine the optimal positions for the virtual UAVs (vUAVs) and the most suitable subchannels as well as power resources among the slices. Then, intra-slice resource sharing is arranged by virtual base stations, vUAVs, or virtual LEO satellites, depending on the distributed units. Eventually, near-Pareto optimal solutions can be found. Simulation results show that, compared to a baseline approach, the proposed framework can improve throughput and delay by up to 10% and 50%, respectively. With a focus on dynamically slicing spectrum resource in SAGINs an online control framework is proposed in [72]. Here, the proposed framework aims to adapt to varying vehicular environments and achieve isolated service provisioning, i.e., each type of service is processed in an independent queue. To that end, the authors propose a workflow of dynamic slicing consisting of four steps, i.e., request admission, request scheduling, UAV dispatching, and resource slicing, to ensure that services with different QoS are adequately served. Based on that, a Lyapunov-based approach is proposed to maximize the system revenue and minimize a time-averaged penalty while stabilizing the system. This approach aims to minimize time-averaged queue backlogs of all services, i.e., the average number of unprocessed requests in a queue over a period. Simulation results show that the proposed framework can increase the throughput by approximately 26%.
@@ III-C1 Resource Allocation
III-C2 Slice Configuration and Routing
Together with resource allocation, multiple aspects are considered to enhance network performance, including slice configuration, Service Level Agreement (SLA) decomposition, and routing. A multi-domain network slicing framework is proposed in [75], aims to jointly optimize all the three aspects. It is shown in [75] that the proposed framework can handle multiple slice configurations in satellite-terrestrial edge computing networks (STECNs). Particularly, a Markov process is modeled to track the probability of satisfying the SLA per configuration. Moreover, an index-based slice configuration policy, based on restless multi-armed bandits (RMABs) [76] [77], is defined to solve the multi-domain slicing problem. Based on the slice configuration policy and the slice/resource availability, the SLA is decomposed. The routing and resource allocation optimization problem is solved by each domain controller with consideration of the slice/resource availability. Simulation results show that the proposed framework can achieve three to six times higher rewards than those of baseline approaches in terms of user requirements in the SLA and the energy consumption. It is also shown in [75] that the proposed framework can achieve optimal configuration under different network conditions, e.g., outage and delay.
@@ III-C2 Slice Configuration and Routing
Summary: In this section, we have discussed multiple approaches for network slicing integration in NTNs, e.g., satellite networks, ISTNs, and SAGINs. Specifically, network slicing is an effective solution to enhance NTNs with better delay, throughput, network resource utilization, network flexibility, and scalability. However, multiple challenges arise when applying network slicing to NTNs such as resource allocation, routing, load balancing, slice admission control, and SLA decomposition. Among those challenges, resource allocation is the most noticeable concern. To address resource allocation, one approach is to formulate an optimization problem, which can be solved by AI/ML-based methods, mathematic algorithms, scheduling strategies, or innovative mechanisms. Additionally, optimizing VNE is effective in enhancing resource allocation. Another challenge to network slicing is routing, which can be solved by shortest-path approaches. Although the proposed approaches can effectively address various challenges of network slicing in NTN, there are still open issues that stem from the complexity of network slicing optimization problems. Particularly, these problems often include numerous constraints due to resource limitations and various platforms, e.g., space, air, and ground, resulting in high complexity. As a result, most of the proposed approaches do not consider network slicing holistically or are based on simplified assumptions. Therefore, additional effort is needed to develop effective and holistic solutions for network slicing in NTNs, especially the integration of NTNs and terrestrial networks. The works surveyed in this section are summarized in Table II.
TABLE II: Summary of network slicing approaches in NTN.
@@ IV-A IoT
In [79], the authors investigate the potential use of DL for NTN-based industrial IoT (IIoT) services, i.e., spaceborne-based and airborne-based intelligent IIoT systems. This research describes how diverse DL-based algorithms have been used and tested for NTN-based IIoT applications based on the optimization objective, the available power, and processing resources. Here, DL is considered as the most effective AI algorithm when it comes to end-to-end optimization. Specifically, for spaceborne applications, an ensemble DNN-based optimization can be utilized for resource allocation while a deep-Q network can be used for energy consumption and processing latency in Satellite-IoT (S-IoT) networks. Additionally, S-IoT edge computing and DL methods can accelerate data processing and transmission as well as enhance bandwidth utilization. Since the practical viability of conventional DL is contingent on the availability of a large amount of sensor data and sufficient computing power, DRL with lower complexity can be used. For airborne applications, an IoT network with three layers for online large data processing based on MEC is proposed with UAVs as edge servers, Lyapunov optimization, DRL algorithm, and CNN Q-networks. In this case, the CNN Q-network, which enables action reward prediction, is trained using the UAVs’ views of the surrounding environment, and the DRL can efficiently optimize the path planning of the UAVs.
@@ IV-B Mobile Services
Since NTN consists of space-borne base stations (BSs), e.g., satellites, and airborne BSs, e.g., UAVs and aircraft, they may lead to a dynamic and non-stationary environment. As such, non-terrestrial users (NT-users) on the ground need to predict the mobility of those non-terrestrial BSs (NT-BSs) autonomously. Likewise, NT-users may move frequently, e.g., drones, which make the trajectory estimation from the NT-BSs side challenging (as illustrated in Fig. 13).
@@ IV-C Network Management
The authors in [84] propose a resource management framework in heterogeneous satellite networks. In the framework, SDN and virtualization methods in the data center are used to manage and combine disparate resources as shown in Fig. 14. To obtain the optimal resource utilization, a DRL approach that uses Markov decision process (MDP) and integrates DL for inference capability as well as RL for decision-making potential is then applied. In this work, the state space contains service and resource states, whereas the action state includes all actions in which an agent provides resources to the services. Additionally, the reward may include spectral efficiency, bandwidth, throughput, and power efficiency. Here, each satellite acts as a smart multi-agent that can perform distributed data processing and transmit information between satellites and/or ground stations in the cloud system. This reduces the workload of the data center and enhances the efficiency of communication because of shorter transmission paths. The SDN-based integrated networks for resource allocation are also proposed in [85]. The purpose of this integrated network is to monitor networks, computation resources as well as caches, and orchestrate them all simultaneously. They use Markov decision process for the resource allocation optimization problem, and deep Q-learning approach for the problem solving. The simulation results reveal that the proposed scheme can achieve the highest expected utility per resource up to 10 times compared with other baseline methods.
@@
Then, the work in [86] presents a novel framework of "self-evolving networks (SENs)," which employs AI through ML algorithms, e.g., federated learning and online learning, to fully automate and intelligently evolve future integrated NTN in terms of network management, communication, computation, and mobility of mobile users. For that, the authors utilize the intelligent vertical heterogeneous network (I-VHetNet) architecture as a model to envision the idea of SEN in future integrated networks. The I-VHetNet design not only combines terrestrial, aerial, and satellite networks, but also includes intelligence, computation, and caching platforms to allow multi-level edge computing. In particular, the SEN engine first employs AI to forecast where additional network capacity and coverage are required based on user movement, behavior, and applications. Utilizing the prediction result, the SEN engine intelligently and automatically sends UAV-BSs or adjusts a HAPS beam to enhance network coverage and increase its capacity to serve customers. It then can choose terrestrial, aerial, or satellite networks to backhaul UAV-BSs based on the QoS requirements. The SEN engine can also execute computational offloading to the best computational level systems, e.g., cloud computing, fog computing, or collaborative computing of mobile users. Afterward, the SEN engine can monitor network performance and user satisfaction. To adapt to dynamic changes and create more accurate automated and intelligent decisions, the SEN engine uses network environment assessment as feedback. From the simulation using three data centers and 300/3000 users, the proposed system can minimize the data offloading and computing delay by 0.05 seconds with 10GB-10TB offloaded data per user.
@@ IV-D Vehicular Networks
Then, the authors in [92] foresee 6G convergent terrestrial and non-terrestrial networks of virtual emotion and pandemic prevention from two perspectives that are Red AI for accuracy and Green AI for efficiency. Specifically, the Red AI-enabled 6G virtual emotion approach leveraging DL algorithms can be used to detect specific emotions of humans passing through a specified area using vehicles with high accuracy and low delay. Furthermore, the Red AI-enabled 6G epidemic prevention approach is utilized to provide epidemic services including fast medical item delivery and epidemic prevention map construction using autonomous vehicles with smart devices on the ground. Meanwhile, Green AI-enabled 6G virtual emotion focuses on computation cost reduction in detecting emotion with pre-defined accuracy. For Green AI-enabled 6G epidemic prevention using DL algorithms, it provides the same services as the Red AI however with the minimum data and reduced number of training processes and communications, thereby maximizing the efficiency. All the above are supported by the NTN via 6G communications.
@@ IV-E UAV
In [95], UAV-based low altitude platform systems (LAPS) with ML/DL approaches to create a flying ad hoc network (FANET) is discussed. Specifically, Q-learning-based RL can be used to develop autonomous and adaptive packet routing among UAVs. In addition to Q-learning, SVM and logistic regression can support real-time and dynamic resource allocation to provide robust UAV services. The combination of Q-learning and FL, i.e., federated Q-learning can also be utilized to protect the FANET through jamming detection.
@@ IV-F Maritime Services
[…]
IV-G Other Applications
@@
Summary: In this section, we have discussed the ML/AI-aided NTN for various emerging applications/services, e.g, IoT systems, mobile services, network management, vehicular networks, UAVs, maritime services, cybertwin, and adaptive data transmission. Particularly, ML/AI approaches have emerged as one of the most potential solutions in dynamic and mobile environments where the NTN radio access network, e.g., satellites, frequently move. Additionally, ML/AI-based NTN can support integrated networks where the deployments of satellites on multiple orbits for space networks, UAVs/drones for aerial networks, and base stations for ground networks exist in such a highly dynamic environment. Here, we can observe from the above discussion that DL, RL, FL, and the combination among them can take part successfully to help NTN implementation for various services. First, the DL method is popular for end-to-end optimization, e.g., resource allocation, channel estimation, and scheduling. In this case, DNN with deep Q-network or the joint DL and RL approaches, i.e., DRL, can be utilized to reduce the complexity of the data training processes. Meanwhile, the use of CNN can be used to optimize the routing strategy of data traffic in the integrated networks. All the aforementioned ML/AI techniques adopt the centralized ML approach in which all radio accesses in the integrated network send their data to a cloud server. Alternatively, the FL approach can be used to provide a decentralized ML approach, where satellites, UAVs, ground base station, and mobile users can train their local data individually and then share the train models with the cloud server. Nonetheless, the utilization of ML/AI for NTN also faces several drawbacks including the increasing complexity of the system, the use of huge amounts of data, and longer training duration. Hence, future research to minimize these problems should be investigated. The summary of all ML/AI-aided NTN for various applications is presented in Table III.
TABLE III: Summary of ML/AI-aided NTN for Various Emerging Applications
@@
IoT
Resource allocation, energy consumption, trajectory optimization
DNN, DRL, CNN
[79]
Power and channel allocation, latency minimization
DRL
[80]
Routing strategy optimization, spectrum sharing
CNN, DQN, SVM
[81]
Mobile services
Handover and throughput optimization
DRL
[82]
Trajectory forecast
SR distributive Q-learning
[83]
Network management
Resource utilization and allocation
DRL
[84, 85]
Network capacity and coverage prediction
FL, online learning
[86]
Network slicing
Transfer and meta learning
[87]
Network slicing and resource allocation
WR-NN
[88]
Resource allocation
FL
[89]
Vehicular networks
Congestion control
DRL, CNN
[90]
Resource slicing and scheduling
RL
[91]
Virtual emotion and pandemic prevention
DL
[92]
UAVs
Antenna tilt deployment optimization
Bi-DNN
[94]
Autonomous and adaptive packet routing, dynamic resource allocation, jamming detection
RL, SVM and logistic regression, Federated RL
[95]
Packet forwarding
DRL
[96]
Maritime services
Energy-efficient topology management and scheduling
DNN, LSTM
[97]
Other applications
Cybertwin-based image classification
FL
[98]
Adaptive data transmission
DRL
[99]
Optimal routing path
CNN
[100]
Clustering-based scheduling problem
K-means, hierarchical clustering, and self-organization
[101]
V ORAN-AIDED NTN

### v2

@@ 3.2.3 Routing
3.3 Space-Air-Ground Integrated Networks
Space-Air-Ground Integrated Networks (SAGINs) aim to achieve full network coverage and ubiquitous services by integrating terrestrial networks, satellite networks, and aerial networks[92, 93]. While applying network slicing to SAGINs is an effective solution for efficient usage of network resources, this technique faces new challenges such as complex slice orchestration in multi-domain networks, resource optimization considering UAVs position, and dispatching cost. To address those challenges, multiple innovative schemes have been proposed.
3.3.1 Resource Allocation
For instance, as the main concern of network slicing application, resource allocation is discussed in [94, 95, 96, 97]. Particularly, in [94], the authors propose a framework for integrating network slicing to a SAGIN, which establishes three types of RAN slices, i.e., high-throughput, low-delay, and wide-coverage. In particular, a non-scalar multi-objective optimization problem (MOOP) is formulated to jointly optimize throughput, service delay, and coverage area. Moreover, a Central and Distributed Multi-agent Deep Deterministic Policy Gradient (CDMADDPG) algorithm is developed to solve the problem. This CDMADDPG algorithm uses a centralized unit to determine the optimal positions for the virtual UAVs (vUAVs) and the most suitable subchannels as well as power resources among the slices. Then, intra-slice resource sharing is arranged by virtual base stations, vUAVs, or virtual LEO satellites, depending on the distributed units. Eventually, near-Pareto optimal solutions can be found. Simulation results show that, compared to a baseline approach, the proposed framework can improve throughput and delay by up to 10% and 50%, respectively. With a focus on dynamically slicing spectrum resource in SAGINs an online control framework is proposed in [95]. Here, the proposed framework aims to adapt to varying vehicular environments and achieve isolated service provisioning, i.e., each type of service is processed in an independent queue. To that end, the authors propose a workflow of dynamic slicing consisting of four steps, i.e., request admission, request scheduling, UAV dispatching, and resource slicing, to ensure that services with different QoS are adequately served. Based on that, a Lyapunov-based approach is proposed to maximize the system revenue and minimize a time-averaged penalty while stabilizing the system. This approach aims to minimize time-averaged queue backlogs of all services, i.e., the average number of unprocessed requests in a queue over a period. Simulation results show that the proposed framework can increase the throughput by approximately 26%.
@@ 3.3.1 Resource Allocation
3.3.2 Slice Configuration and Routing
Together with resource allocation, multiple aspects are considered to enhance network performance, including slice configuration, Service Level Agreement (SLA) decomposition, and routing. A multi-domain network slicing framework is proposed in [98], which aims to jointly optimize all the three aspects. It is shown in [98] that the proposed framework can handle multiple slice configurations in satellite-terrestrial edge computing networks (STECNs). Particularly, a Markov process is modeled to track the probability of satisfying the SLA per configuration. Moreover, an index-based slice configuration policy, based on restless multi-armed bandits (RMABs) [99] [100], is defined to solve the multi-domain slicing problem. Based on the slice configuration policy and the slice/resource availability, the SLA is decomposed. The routing and resource allocation optimization problem is solved by each domain controller with consideration of the slice/resource availability. Simulation results show that the proposed framework can achieve three to six times higher rewards than those of baseline approaches in terms of user requirements in the SLA and the energy consumption. It is also shown in [98] that the proposed framework can achieve optimal configuration under different network conditions, e.g., outage and delay.
@@ 3.3.2 Slice Configuration and Routing
Summary: In this section, we have discussed multiple approaches for network slicing integration in NTNs, e.g., satellite networks, ISTNs, and SAGINs. Specifically, network slicing is an effective solution to enhance NTNs with better delay, throughput, network resource utilization, network flexibility, and scalability. However, multiple challenges arise when applying network slicing to NTNs such as resource allocation, routing, load balancing, slice admission control, and SLA decomposition. Among those challenges, resource allocation is the most noticeable concern. To address resource allocation, one approach is to formulate an optimization problem, which can be solved by AI/ML-based methods, mathematic algorithms, scheduling strategies, or innovative mechanisms. Additionally, optimizing VNE is effective in enhancing resource allocation. Another challenge to network slicing is routing, which can be solved by shortest-path approaches. Although the proposed approaches can effectively address various challenges of network slicing in NTN, there are still open issues that stem from the complexity of network slicing optimization problems. Particularly, these problems often include numerous constraints due to resource limitations and various platforms, e.g., space, air, and ground, resulting in high complexity. As a result, most of the proposed approaches do not consider network slicing holistically or are based on simplified assumptions. Therefore, additional effort is needed to develop effective and holistic solutions for network slicing in NTNs, especially the integration of NTNs and terrestrial networks. The works surveyed in this section are summarized in Table 3.
Table 3: Summary of network slicing approaches in NTN.
@@ 4.1 IoT
In [102], the authors investigate the potential use of DL for NTN-based industrial IoT (IIoT) services, i.e., spaceborne-based and airborne-based intelligent IIoT systems. This research describes how diverse DL-based algorithms have been used and tested for NTN-based IIoT applications based on the optimization objective, the available power, and processing resources. Here, DL is considered the most effective AI algorithm when it comes to end-to-end optimization. Specifically, for spaceborne applications, an ensemble DNN-based optimization can be utilized for resource allocation while a deep-Q network can be used for energy consumption and processing latency in Satellite-IoT (S-IoT) networks. Additionally, S-IoT edge computing and DL methods can accelerate data processing and transmission as well as enhance bandwidth utilization. Since the practical viability of conventional DL is contingent on the availability of a large amount of sensor data and sufficient computing power, DRL with lower complexity can be used. For airborne applications, an IoT network with three layers for online large data processing based on MEC is proposed with UAVs as edge servers, Lyapunov optimization, DRL algorithm, and CNN Q-networks. In this case, the CNN Q-network, which enables action reward prediction, is trained using the UAVs’ views of the surrounding environment, and the DRL can efficiently optimize the path planning of the UAVs.
@@ 4.2 Mobile Services
Since NTN consists of space-borne base stations (BSs), e.g., satellites, and airborne BSs, e.g., UAVs and aircraft, they may lead to a dynamic and non-stationary environment. As such, non-terrestrial users (NT-users) on the ground need to predict the mobility of those non-terrestrial BSs (NT-BSs) autonomously. Likewise, NT-users may move frequently, e.g., drones, which makes the trajectory estimation from the NT-BSs side challenging (as illustrated in Fig. 13).
@@ 4.3 Network Management
The authors in [107] propose a resource management framework in heterogeneous satellite networks. In the framework, SDN and virtualization methods in the data center are used to manage and combine disparate resources as shown in Fig. 14. To obtain the optimal resource utilization, a DRL approach that uses the Markov decision process (MDP) and integrates DL for inference capability as well as RL for decision-making potential is then applied. In this work, the state space contains service and resource states, whereas the action state includes all actions in which an agent provides resources to the services. Additionally, the reward may include spectral efficiency, bandwidth, throughput, and power efficiency. Here, each satellite acts as a smart multi-agent that can perform distributed data processing and transmit information between satellites and/or ground stations in the cloud system. This reduces the workload of the data center and enhances the efficiency of communication because of shorter transmission paths. The SDN-based integrated networks for resource allocation are also proposed in [108]. The purpose of this integrated network is to monitor networks, computation resources as well as caches, and orchestrate them all simultaneously. They use the Markov decision process for the resource allocation optimization problem, and deep Q-learning approach for the problem solving. The simulation results reveal that the proposed scheme can achieve the highest expected utility per resource up to 10 times compared with other baseline methods.
@@
Then, the work in [109] presents a novel framework of ”self-evolving networks (SENs),” which employs AI through ML algorithms, e.g., federated learning and online learning, to fully automate and intelligently evolve future integrated NTN in terms of network management, communication, computation, and mobility of mobile users. For that, the authors utilize the intelligent vertical heterogeneous network (I-VHetNet) architecture as a model to envision the idea of SEN in future integrated networks. The I-VHetNet design not only combines terrestrial, aerial, and satellite networks, but also includes intelligence, computation, and caching platforms to allow multi-level edge computing. In particular, the SEN engine first employs AI to forecast where additional network capacity and coverage are required based on user movement, behavior, and applications. Utilizing the prediction result, the SEN engine intelligently and automatically sends UAV-BSs or adjusts a HAPS beam to enhance network coverage and increase its capacity to serve customers. It then can choose terrestrial, aerial, or satellite networks to backhaul UAV-BSs based on the QoS requirements. The SEN engine can also execute computational offloading to the best computational level systems, e.g., cloud computing, fog computing, or collaborative computing of mobile users. Afterward, the SEN engine can monitor network performance and user satisfaction. To adapt to dynamic changes and create more accurate automated and intelligent decisions, the SEN engine uses network environment assessment as feedback. From the simulation using three data centers and 300/3000 users, the proposed system can minimize the data offloading and computing delay by 0.05 seconds with 10GB-10TB offloaded data per user.
@@ 4.4 Vehicular Networks
Then, the authors in [115] foresee 6G convergent terrestrial and NTNs of virtual emotion and pandemic prevention from two perspectives that are Red AI for accuracy and Green AI for efficiency. Specifically, the Red AI-enabled 6G virtual emotion approach leveraging DL algorithms can be used to detect specific emotions of humans passing through a specified area using vehicles with high accuracy and low delay. Furthermore, the Red AI-enabled 6G epidemic prevention approach is utilized to provide epidemic services including fast medical item delivery and epidemic prevention map construction using autonomous vehicles with smart devices on the ground. Meanwhile, Green AI-enabled 6G virtual emotion focuses on computation cost reduction in detecting emotion with pre-defined accuracy. For Green AI-enabled 6G epidemic prevention using DL algorithms, it provides the same services as the Red AI however with the minimum data and reduced number of training processes and communications, thereby maximizing the efficiency. All the above are supported by the NTN via 6G communications.
@@ 4.5 UAV
In [118], UAV-based low altitude platform systems (LAPS) with ML/DL approaches to create a flying ad hoc network (FANET) is discussed. Specifically, Q-learning-based RL can be used to develop autonomous and adaptive packet routing among UAVs. In addition to Q-learning, SVM and logistic regression can support real-time and dynamic resource allocation to provide robust UAV services. The combination of Q-learning and FL, i.e., federated Q-learning can also be utilized to protect the FANET through jamming detection. A case study with ground base station and UAVs in smart farming scenarios using the opportunistic network environment shows that the proposed system can achieve more than 95% delivery ratio and average latency 50% lower than those of other delay tolerant networks.
@@ 4.6 Maritime Services
For example, the use of ML/AI-based methods to provide sustainable 6G maritime networks through NTN is investigated in [120]. Particularly, the most potential maritime scenarios are first described including maritime search and rescue, intelligent harbor and vessel logistics, on-board entertainment, navigation and fleet management, and shipborne IoT. In this case, distributed intelligence, wherein learning and inference are utilized at several system levels, is required to cope with the dynamic networking. Then, the authors show how ML may improve energy-efficient topology management and scheduling in dynamic marine networks over baseline model-based techniques. For the energy-efficient topology management, a multi-hop wireless network with the aid of NTN is deployed in the area of 100 km2 to reduce energy usage of the network while finding the optimal routing paths for heterogeneous traffic delivery to multiple destinations, e.g., cargo vessels. For that, the DNN-based DL approach is utilized to estimate which connections are absolutely necessary for the best possible configuration as the system’s traffic patterns change. As a result, the number of involving connections in the routing and execution time can be minimized. Additionally, the temporal and geographical correlations in traffic may be discovered and multi-objective optimization issues can be supported by employing other promising methods such as LSTM and auto-encoder. For the energy-efficient scheduling, the LSTM approach can be applied to predict channel quality because of its ability to deal with time series problems. Consequently, the channel reporting accuracy can be improved and packet delay can be reduced at a fixed 2 seconds for a user density of more than 2 users/m2.
4.7 Other Applications
@@
Summary: In this section, we have discussed the ML/AI-aided NTN for various emerging applications/services, e.g, IoT systems, mobile services, network management, vehicular networks, UAVs, maritime services, cybertwin, and adaptive data transmission. Particularly, ML/AI approaches have emerged as one of the most potential solutions in dynamic and mobile environments where the NTN radio access network, e.g., satellites, frequently move. Additionally, ML/AI-based NTN can support integrated networks where the deployments of satellites on multiple orbits for space networks, UAVs/drones for aerial networks, and base stations for ground networks exist in such a highly dynamic environment. Here, we can observe from the above discussion that DL, RL, FL, and the combination among them can take part successfully to help NTN implementation for various services. First, the DL method is popular for end-to-end optimization, e.g., resource allocation, channel estimation, and scheduling. In this case, DNN with deep Q-network or the joint DL and RL approaches, i.e., DRL, can be utilized to reduce the complexity of the data training processes. Meanwhile, the use of CNN can be used to optimize the routing strategy of data traffic in the integrated networks. All the aforementioned ML/AI techniques adopt the centralized ML approach in which all radio accesses in the integrated network send their data to a cloud server. Alternatively, the FL approach can be used to provide a decentralized ML approach, where satellites, UAVs, ground base station, and mobile users can train their local data individually and then share the train models with the cloud server. Nonetheless, the existing works of ML/AI for NTN also face several drawbacks. First, ML/AI-aided NTN may increase the complexity of the overall NTN system due to the integration of various ML/AI methods and their compatibility with existing network infrastructure. Here, the next research can focus on simplifying the implementation process without compromising accuracy performance through modular design, protocol standardization, and abstraction layer introduction of ML/AI method in NTN. Second, NTN with ML/AI method can also suffer from large amounts of data that can lead to data storage, processing, and transmission problems. To address this issue, future research can explore techniques for data compression, feature selection, or data augmentation in NTN. Third, the ML/AI-based NTN may introduce computationally-intensive and time-consuming training process which can delay the deployment of NTN solutions. To this end, future research can focus on developing more efficient training algorithms, optimizing hardware accelerators, or exploring distributed training approaches to parallelize the training process. Finally, based on the existing centralized and decentralized ML/AI approaches, future research can investigate the trade-offs between these approaches in terms of privacy, latency, and scalability, and then develop hybrid approaches that combine the benefits of both centralized and decentralized ML/AI in NTN. The summary of all ML/AI-aided NTN for various applications is presented in Table 4.
Table 4: Summary of ML/AI-aided NTN for Various Emerging Applications
@@
IoT Resource allocation, energy consumption, trajectory optimization DNN, DRL, CNN [102]
Power and channel allocation, latency minimization DRL [103]
Routing strategy optimization, spectrum sharing CNN, DQN, SVM [104]
Mobile services Handover and throughput optimization DRL [105]
Trajectory forecast SR distributive Q-learning [106]
Network management Resource utilization and allocation DRL [107, 108]
Network capacity and coverage prediction FL, online learning [109]
Network slicing Transfer and meta learning [110]
Network slicing and resource allocation WR-NN [111]
Resource allocation FL [112]
Vehicular networks Congestion control DRL, CNN [113]
Resource slicing and scheduling RL [114]
Virtual emotion and pandemic prevention DL [115]
UAVs Antenna tilt deployment optimization Bi-DNN [117]
Autonomous and adaptive packet routing, dynamic resource allocation, jamming detection RL, SVM and logistic regression, Federated RL [118]
Packet forwarding DRL [119]
Maritime services Energy-efficient topology management and scheduling DNN, LSTM [120]
Other applications Cybertwin-based image classification FL [121]
Adaptive data transmission DRL [122]
Optimal routing path CNN [123]
Clustering-based scheduling problem K-means, hierarchical clustering, and self-organization [124]
5 ORAN-aided NTN